simplify and write your answer with a positive exponent.
x to the -8 power OVER x to the -12 power

Answer by scott8148(6628)

when dividing exponentiated quantities that have the same base, subtract the exponents

(x^(-8))/(x^(-12))=x^(-8-(-12))=x^4